### v2

@@ B.7 Peer Influence pre and post 1850
Share Free-State peers (sd) 0.041 0.034 0.044 0.048
@@
Dependent var.mean 0.391 0.391 0.318 0.318
Observations 235 235 154 154
R-squared 0.227 0.281 0.220 0.333
@@ B.8 Controlling for Five-Year Fixed Effects
Share Free-State peers (sd) 0.061∗∗ 0.048∗∗ 0.043 0.002 0.002 -0.001
@@
Dependent var. mean 0.362 0.362 0.362 0.919 0.919 0.919
Observations 389 389 389 540 540 540
R-squared 0.031 0.225 0.272 0.033 0.041 0.055
Note. This table presents the impact of the fraction of peers from Free States in a cadet’s cohort on that cadet’s decision to join the Union. We include fixed effects at five year intervals to allow for a shift in the environment. Columns (1)-(3) focus on cadets from Slave States and Columns (4)-(6) on those from Free States. The standard errors presented in the parentheses are obtained through bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.
@@
Share Free-State total peers (sd) 0.043 0.039
@@
Share Free-State graduates (sd) 0.044 0.039
@@
Share Free-State dropouts (sd) 0.001 0.001
@@
Dependent var.mean 0.362 0.362 0.362 0.362
Observations 389 389 389 389
R-squared 0.219 0.266 0.220 0.267
@@ B.10 Multinomial Logit Results
Share Free-State peers (sd): Join the Union 0.267∗∗ 0.261∗∗ 0.269∗∗
@@
Share Free-State peers (sd): Not in war 0.044 0.040 0.055
@@
Dependent var. mean 1.214 1.214 1.214
Observations 674 674 674
Pseudo R-squared 0.005 0.179 0.203
Note. The sample for this analysis consists of all cadets from Slave States, including those who did not join either army. The results show that peer influence mainly affected whether to join the Union or Confederacy rather than whether to engage in the war. The controls are the same as in our previous analyses. The standard errors reported in the parentheses are derived from bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.
@@ B.11 Considering Cash-Crop Employment
(1) (2) (3) (4) (5) (6)
Share Free-State peers (sd) 0.062∗∗ 0.044∗∗ 0.039 0.006 -0.000 -0.003
@@
Age in 1860 -0.014 -0.003 0.007 0.005
(0.015) (0.016) (0.007) (0.007)
Class rank 0.001 0.001 -0.001 -0.001
(0.001) (0.001) (0.000) (0.000)
Slave pop. share (sd) -0.201∗∗∗ -0.005
@@
Cash crop employment -0.264∗∗∗ -0.247∗∗∗ -0.624∗∗∗ -0.635∗∗∗
(0.054) (0.052) (0.187) (0.184)
Cohort -0.016 -0.005 0.008 0.007
(0.015) (0.015) (0.006) (0.006)
State FEs N N Y N N Y
Dependent var. mean 0.362 0.362 0.362 0.919 0.919 0.919
Observations 389 389 389 540 540 540
R-squared 0.017 0.236 0.282 0.000 0.081 0.096
@@
Table S15: Peer composition and allegiance choice: interaction term of cash crops
Dependent var Join the Union: war participants
Slave States Free States
(1) (2) (3) (4)
Cash crop $\times$ Share Free-State peers (sd) 0.055 0.036 0.022 0.032
(0.057) (0.062) (0.709) (0.683)
Share Free-State peers (sd) 0.042 0.037 -0.000 -0.003
(0.023) (0.024) (0.012) (0.013)
Age in 1860 -0.015 -0.003 0.007 0.005
(0.016) (0.016) (0.007) (0.007)
Class rank 0.001 0.001 -0.001 -0.001
(0.001) (0.001) (0.000) (0.000)
Slave pop. share (sd) -0.203∗∗∗ -0.005
(0.019) (0.013)
Cash crop employment -0.253∗∗∗ -0.240∗∗∗ -0.613 -0.618
(0.056) (0.053) (0.670) (0.649)
Cohort -0.017 -0.005 0.008 0.006
(0.015) (0.015) (0.006) (0.006)
State FEs N Y N Y
Dependent var. mean 0.362 0.362 0.919 0.919
Observations 389 389 540 540
R-squared 0.237 0.282 0.081 0.096
Note. This table presents the impact of the fraction of peers from Free States in a cadet’s cohort on that cadet’s decision to join the Union. Cash crops is a dummy variable that equals 1 when Graduates with connections to Southern cash crops. Columns (1)-(2) focus on cadets from Slave States and Columns (3)-(4) on those from Free States. The standard errors presented in the parentheses are obtained through bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.
@@ B.12 Considering Appointment State Peers
Share Free-State peers (sd) 0.056∗∗ 0.046∗∗ 0.039 -0.003 -0.006 -0.008
@@
Dependent var. mean 0.362 0.362 0.362 0.919 0.919 0.919
Observations 389 389 389 540 540 540
R-squared 0.014 0.220 0.268 0.000 0.016 0.029
Note. This table presents the impact of the fraction of peers from Free States in a cadet’s cohort on that cadet’s decision to join the Union. Share Free-State peers (sd) of appointed state is the proportion of free companions calculated based on the place of appointment of graduates. Columns (1)-(3) focus on cadets from Slave States and Columns (4)-(6) on those from Free States. The standard errors presented in the parentheses are obtained through bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.

### v3

@@ B.7 Peer Influence pre and post 1850
Share Free-State Peers (sd) 0.054∗∗ 0.045 0.057 0.062
@@
Dependent var.mean 0.389 0.389 0.318 0.318
Observations 234 234 154 154
R-squared 0.236 0.301 0.231 0.332
@@ B.8 Controlling for Five-Year Fixed Effects
Share Free-State Peers (sd) 0.076∗∗∗ 0.062∗∗∗ 0.058∗∗ -0.008 -0.009 -0.011
@@
Dependent var. mean 0.361 0.361 0.361 0.919 0.919 0.919
Observations 388 388 388 540 540 540
R-squared 0.041 0.240 0.292 0.033 0.043 0.057
Note. This table presents the impact of the fraction of peers from Free States in a cadet’s cohort on that cadet’s decision to join the Union. We include fixed effects at five year intervals to allow for a shift in the environment. The control variables consist of Age in 1860, Class Rank, state level Slave Population Share (sd), and Cohort. Columns (1)-(3) focus on cadets from Slave States and Columns (4)-(6) on those from Free States. The standard errors presented in the parentheses are obtained through bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.
@@
Share Free-State Total Peers (sd) 0.058∗∗ 0.055∗∗
@@
Share Free-State Graduates (sd) 0.059∗∗∗ 0.053∗∗
@@
Share Free-State Dropouts (sd) 0.009 0.010
@@
Dependent var.mean 0.361 0.361 0.361 0.361
Observations 388 388 388 388
R-squared 0.231 0.286 0.233 0.286
@@ B.10 Multinomial Logit Results
Share Free-State Peers (sd): Join the Union 0.349∗∗∗ 0.336∗∗∗ 0.358∗∗∗
@@
Share Free-State Peers (sd): Not in war 0.102 0.063 0.079
@@
Dependent var. mean 1.212 1.212 1.212
Observations 670 670 670
Pseudo R-squared 0.008 0.181 0.206
Note. The sample for this analysis consists of all cadets from Slave States, including those who did not join either army. The results show that peer influence mainly affected whether to join the Union or Confederacy rather than whether to engage in the war. The control variables consist of Age in 1860, Class Rank, state level Slave Population Share (sd), and Cohort. The standard errors reported in the parentheses are derived from bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.
@@ B.11 Considering Cash-Crop Employment
(1) (2) (3) (4) (5) (6) (7) (8)
Share Free-State Peers (sd) 0.080∗∗∗ 0.058∗∗∗ 0.053∗∗∗ 0.052∗∗ -0.003 -0.009 -0.011 -0.011
(0.023) (0.020) (0.020) (0.020) (0.013) (0.012) (0.012) (0.012)
Age in 1860 -0.013 -0.006 -0.006 0.006 0.004 0.003
@@
Class rank 0.001 0.001 0.001 -0.001 -0.001 -0.001
(0.001) (0.001) (0.001) (0.000) (0.000) (0.000)
Slave Pop. Share (sd) -0.203∗∗∗ -0.004
@@
Cash Crop Employment -0.259∗∗∗ -0.234∗∗∗ -0.230∗∗∗ -0.627∗∗∗ -0.638∗∗∗ -0.621
(0.056) (0.054) (0.054) (0.187) (0.187) (3.384)
Cash Crop $\times$ Share Free-State Peers (sd) 0.027 0.033
(0.066) (4.332)
Cohort -0.014 -0.007 -0.007 0.007 0.006 0.005
(0.015) (0.015) (0.015) (0.007) (0.007) (0.007)
State FEs N N Y Y N N Y Y
Dependent var. mean 0.361 0.361 0.361 0.361 0.919 0.919 0.919 0.919
Observations 388 388 388 388 540 540 540 540
R-squared 0.028 0.250 0.300 0.300 0.000 0.084 0.098 0.099
@@ B.12 Considering Appointment State Peers
Share Free-State peers (sd) 0.066∗∗∗ 0.052∗∗∗ 0.046∗∗ -0.009 -0.012 -0.014
@@
Dependent var. mean 0.361 0.361 0.361 0.919 0.919 0.919
Observations 388 388 388 540 540 540
R-squared 0.019 0.231 0.284 0.001 0.018 0.032
Note. This table presents the impact of the fraction of peers from Free States in a cadet’s cohort on that cadet’s decision to join the Union. Share Free-State peers (sd) of appointed state is the proportion of free companions calculated based on the place of appointment of graduates. The control variables consist of Age in 1860, Class Rank, state level Slave Population Share (sd), and Cohort. Columns (1)-(3) focus on cadets from Slave States and Columns (4)-(6) on those from Free States. The standard errors presented in the parentheses are obtained through bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.
B.13 Career Outcomes
Table S17: Subsequent outcomes in 1865
OLS IV
Rank General Died Rank General Died
(1) (2) (3) (4) (5) (6)
A. Slave-State Cadets (>1%)
Joining the Union -1.567∗∗∗ -0.352∗∗∗ -0.166∗∗∗ -1.858∗∗∗ -0.365∗∗∗ -0.135
(0.194) (0.054) (0.046) (0.346) (0.099) (0.071)
Controls Y Y Y Y Y Y
State FEs Y Y Y Y Y Y
Dependent var. mean 6.987 0.438 0.206 6.987 0.438 0.206
Observations 381 381 388 381 381 388
F-statistic 9.911 9.911 10.255
R-squared 0.317 0.190 0.097 0.267 0.149 0.058
B. Slave-State Cadets (1%-33%)
Joining the Union -1.505∗∗∗ -0.314∗∗∗ -0.170∗∗∗ -1.717∗∗∗ -0.365∗∗ -0.168
(0.223) (0.061) (0.048) (0.513) (0.147) (0.091)
Controls Y Y Y Y Y Y
State FEs Y Y Y Y Y Y
Dependent var. mean 6.802 0.407 0.184 6.802 0.407 0.184
Observations 263 263 266 263 263 266
F-statistic 8.916 8.916 9.189
R-squared 0.314 0.191 0.082 0.262 0.144 0.070
Note. The sample for Panel A consists of the war-participating cadets from Slave States (i.e., with a slave population higher than 1%), and the sample for Panel B excludes those from the High-Slave States who were not significantly influenced by peers. The control variables consist of Age in 1860, Class Rank, and Cohort. Columns (1)-(3) use OLS estimation, while Columns (4)–(6) present IV estimates using birthplace indicators and the share of peers from Free States as instruments for Union affiliation. Columns (1) and (4) feature the military rank as the dependent variable, classified into 11 levels, with General assigned 11 and Third Lieutenant assigned 1. General and Died are binary variables. The rank information on seven cadets is unknown. The standard errors reported in the tables are derived from bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.
B.14 Correlation between Academic Rank and Military Rank in 1865
Table S18: Academic Rank and Military Rank in 1865 (The Confederacy vs. The Union)
Dependent var Rank in 1865: War Participants
The Confederacy The Union
(1) (2) (3) (4) (5) (6)
Class Rank 0.000 -0.001 -0.001 0.007∗∗ 0.007∗∗∗ 0.006∗∗
(0.003) (0.003) (0.003) (0.003) (0.003) (0.003)
Age in 1860 -0.008 -0.027 -0.031 -0.037
(0.052) (0.059) (0.042) (0.046)
Slave Pop. Share (sd) 0.157 0.033
(0.089) (0.064)
Cohort -0.053 -0.074 -0.085∗∗ -0.092∗∗
(0.050) (0.057) (0.040) (0.044)
State FEs N N Y N N Y
Dependent var.mean 7.516 7.516 7.516 5.851 5.851 5.851
Observations 281 281 281 636 636 636
R-squared 0.000 0.086 0.163 0.010 0.111 0.129